October 23th
------------

The retroreflecting optics has been mounted on the telescope and removed around 22:00 when the weather conditions were improved.

Daytime activity
-----------------

1) AO system calibration and performance

-The measurements of the 2 reconstructors at BIN1 have been done. The new reconstructors have modulation 1*Lambda/D and 0.5*lambda/D.
-Interaction matrix of BIN2 recosntructor has been acquired at different clocking to verify the goodness of the assumed clocking for the nominal BIN2 reconstructor. The data confirmed that the chosen position has the better performance of those measured.

Night time activity
---------------------

The observations has started at 23.00 due to wetaher conditions. Two hours has been devoteed to the RFBG focal station test. After this the AO observations started.

AO closed loop test. The first AO star has been pointed at 01:40. The seeing conditions have been not good with oscillations between 0.9 and 1.5. The loop has been succesfully closed using AOS achieving SR in H band between 86 and 56. The SR values correltes with the instantaneous seeing value and are consistent with our numerocal simulations of SR vs. seeing value. Data reduction to correlate SR value with seeing value are ongoing.

AOS test. Because the seeing conditions were worsening (seeing value around or more than 2 arcesc) we decided to perform some SW test with AOS. The test are aimed to demonstrate the automatic setup and loop closure using the high level SW (AOS).
All the AOS commands has been tested succesully. A list of teh command is given below

- preset_ao
- acquire_ref
- check_ref
- start_ao
- offset_x
- offset_y
- offset_z
- pause
- resume
- stop

The time required to issue the three comnads that configure the system, acquire the star and close the loop (preset_ao, acquire_ref, start_ao) has been measured to be in between 2.5 and 3.0 minutes.

Test with the irc commands

The test was done using the BIN2 of the WFS. the following commands has been succesfully tested.

- preset_ao
- acquire_ref
- start_ao
- offset_x
- offset_y
- offset_z
- pause
- resume
- stop


A problem was found in the TCS/IIF command preset_telescope in adaptive ACE mode did not allows to close the loop using IIF commands. A patch for this problem has been developed by the LBTO SW group and we are planning to repeat the test with the upgraded SW.
Topic revision: r1 - 29 Oct 2010, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback